OpenStudy (anonymous):

how to put 2y=12x+7 into general form. I don't remember how to. Thanks

OpenStudy (anonymous):

y=(12x+7)/2

OpenStudy (anonymous):

but itsn't the genral form f(x)=ax+by+c

OpenStudy (anonymous):

you're wanting it in x=?

OpenStudy (asnaseer):

This might help you: http://www.mathsisfun.com/algebra/line-equation-general-form.html

OpenStudy (anonymous):

x=(y/6) - 7/12

OpenStudy (anonymous):

no that is convert to slope-intercept form

OpenStudy (anonymous):

no, cameronmx9, the general form is = to 0 so that mean the whole equation equals to 0

OpenStudy (anonymous):

2y-12x-7=0

OpenStudy (anonymous):

ok that helped, thanks

OpenStudy (anonymous):

Its wrong, but still thanks.

OpenStudy (anonymous):

f(x)=12x-2y+7. What exactly do you want done to the equation? Find slope?? Find Y??

OpenStudy (anonymous):

Or you could say.. -12x+2y+7=0